QRS Corp. is designing a product layout for a new product. They plan to use this production line 10 hours a day in order to meet forecasted demand of 900 units per day. The following table describes the tasks necessary to produce this product:

Task Time(secs) Immediate
Predecessor
a 34 none
b 20 a
c 10 b
d 16 b
e 10 c
f 24 d,e
g 38 f

For output to equal forecasted demand, what will be the second task performed at the second workstation of the production line that uses the fewest number of stations?

a

b

c

d

e
Answer #1

Total Available time = 10 hour

Demand = 900 units /10 hour = 90 units / hr

Cycle time = 60*60/90 = 40 sec.

Total Task time = 152 seconds

Theoretical minimum number of workstations = Total Task time / cycle time = 152 / 40 =3.8 = 4

The work allocation is as follows.

Work Station Tasks Time Idle time
1 a 34 6
2 b,c,e 40 0
3 d,f 40 0
4 g 38 2

From above, The required task is C,